# Look Twice before You Leap: A Rational Agent Framework for Localized Adversarial Anonymization

### v1

@@ 3.3.3 The Anonymizer ($M_{ano}$)
11:   $\pi_{k}\leftarrow\Pi_{select}(v_{k})$ $\triangleright$ Map validity to instruction
@@
4 Experiments
4.1 Experimental Setup
Datasets
Two benchmark datasets are employed to evaluate our approach. PersonalReddit staab24beyond is a dataset widely adopted in researches like IncogniText frikha2024incognitext and RUPTA yang2024robust, consists of 525 human-verified synthetic Reddit conversations annotated with 8 fine-grained private attributes (e.g., age, gender and location). reddit-self-disclosure dou2024reducing serves as a real-world evaluation bed, from which we extracted 885 samples containing self-disclosures of health conditions. Since this dataset is inherently annotated, we utilized the original labels as ground truth and performed manual verification to ensure consistency.
Models
Our experiments involve two distinct model roles: the local model used for the anonymization process and the external evaluation model used to benchmark privacy and utility.
For all local framework components (e.g., the arbitrator in RLAA and the attacker, anonymizer in FgAA (Naive) and SFT migration variant), we tested Llama3-8B and Qwen2.5-7B. All of these models only require about 4GB of video memory to load when using 4-bit quantization, thus enabling it to run locally on most personal computers. The operation of our entire RLAA framework is confined to this local model, structurally eliminating the service provider $\mathcal{A}_{service}$ ’s threat. To evaluate the privacy protection of all methods, we adopt DeepSeek-V3.2-Exp as the unified re-identification attacker, which is a 685B parameter model. We choose DeepSeek-V3.2-Exp for its SoTA performance, which allows us to simulate the most powerful $\mathcal{A}_{re-id}$ (re-identification) adversary.
This evaluation design using a powerful SoTA model to attack outputs generated by a local-scale model is intentional and critical for realistic threat modeling. It precisely simulates the scenario defined in Section 3.1: A user runs RLAA locally to anonymize text, solving the "privacy paradox" ($\mathcal{A}_{service}$). The user then publishes the anonymized text $x^{ano}$, exposing it to external adversaries. A powerful external adversary ($\mathcal{A}_{re-id}$), simulated by DeepSeek-V3.2-Exp, attempts to re-identify the user from $x^{ano}$. This apparent "train-test mismatch" (8B internal attacker vs. 685B external evaluator) is therefore not a flaw, but a crucial stress test to validate whether our fully local framework generates outputs robust enough to withstand worst-case and SoTA-level attacks.
Baseline Methods
We compared RLAA against FgAA staab2024large (including its Naive and SFT migration variants), SEAL kim2025self, IncogniText frikha2024incognitext, and DP-BART-PR+ igamberdiev2023dp. The maximum adversarial iteration limit $T$ was configured according to dataset complexity and the need to evaluate varying convergence horizons: we set $T=10$ for the complex, multi-attribute PersonalReddit and $T=3$ for the simpler, single-attribute reddit-self-disclosure. For DP-BART-PR+, we follow the description in the original paper and set a privacy budget of $\epsilon=2500$. Detailed training configurations for the SFT baseline (e.g., QLoRA settings and learning rate) are provided in Appendix C.
Evaluation Metrics
To evaluate the performance of all methods, we adopt the standardized privacy and utility evaluation protocol widely established in recent LLM-based anonymization research staab2024large; frikha2024incognitext; kim2025self. Let $D=\{(x_{i}^{ori},A_{i})\}_{i=1}^{N}$ be the test dataset, where $x_{i}^{ori}$ is the original text and $A_{i}$ is the set of ground-truth private attributes. Let $x_{i}^{ano}$ be the anonymized text generated by a given method.
Privacy (PRIV)
Privacy is evaluated using the anonymization under LLM inference setting staab2024large. We use a powerful adversary model $\mathcal{M}_{adv}$ (DeepSeek-V3.2-Exp deepseekai2024deepseekv32) to infer the set of attributes $A^{\prime}_{i}=\mathcal{M}_{adv}(x_{i}^{ano})$ from the anonymized text. The privacy score is the average percentage of correct inferences (attack success rate) over all attributes $K$ across all $N$ samples. Note that a lower PRIV score indicates better privacy protection.
$\text{PRIV}=\frac{1}{N\cdot K}\sum_{i=1}^{N}\sum_{k=1}^{K}\mathbb{I}(A^{\prime}_{i,k}=A_{i,k})$
Utility (UTIL)
Utility evaluation relies on an LLM-as-Judge $\mathcal{M}_{judge}$ to assess semantic preservation via $K=3$ distinct components. We denote the scalar score range as $S_{\min},S_{\max}]$ (where $S_{\min}=1,S_{\max}=10$): (1) Readability ($s_{\text{read}}$) is a scalar score $\in[S_{\min},S_{\max}]$ assessing if the text is understandable on its own, defined as $s_{\text{read},i}=\mathcal{M}_{judge}(x_{i}^{ano})$. (2) Meaning ($s_{\text{mean}}$) is a scalar score $\in[S_{\min},S_{\max}]$ assessing if the core message is preserved, where $s_{\text{mean},i}=\mathcal{M}_{judge}(x_{i}^{ori},x_{i}^{ano})$. (3) Hallucinations ($s_{\text{hall}}$) is a binary score $s_{\text{hall},i}=\mathbb{I}(\text{no new ungrounded information})$.
The final composite utility score UTIL is the average of these $K$ normalized components:
Traditional Structural Metrics
In addition to LLM-based utility, we report standard n-gram metrics to evaluate structural similarity, which are calculated separately from the composite UTIL score: ROUGE-L presents The F1 score of the longest common subsequence, and BLEU indicates the bilingual evaluation understudy score for precision.
Dataset: PersonalReddit
Original Text 1.0000 0.4442 1.0000 1.0000 10.000 10.000 1.0000
Local Baselines
DP-BART+ BART-Base 0.3470 0.2650 0.3925 0.2597 2.6610 7.5790 0.0170
FgAA (Naive) Llama3-8B 0.7297 0.1948 0.2180 0.0533 3.5420 8.9330 0.9420
RLAA Llama3-8B 0.8788 0.2130 0.5958 0.4251 7.0780 9.8090 0.9480
API Required Baselines
SEAL (SFT+DPO) Llama3-8B 0.4642 0.1787 0.1205 0.0753 9.2645 1.5207 0.3140
FgAA (SFT) Llama3-8B 0.9670 0.2940 0.9149 0.9389 9.2310 9.9340 0.9830
FgAA (API) DeepSeek-V3.2-Exp 0.8264 0.2056 0.4649 0.2082 5.4380 9.3550 1.0000
Dataset: reddit-self-disclosure
Original Text 1.0000 0.4943 1.0000 1.0000 10.000 10.000 1.0000
Local Baselines
DP-BART+ BART-Base 0.3999 0.3245 0.2565 0.1093 2.3698 8.6830 0.0943
FgAA (Naive) Llama3-8B 0.8187 0.1591 0.4919 0.2805 5.3523 9.6629 0.9545
RLAA Llama3-8B 0.8572 0.1136 0.6016 0.4703 6.8939 9.6932 0.9129
API Required Baselines
SEAL (SFT+DPO) Llama3-8B 0.6303 0.0226 0.1213 0.0736 9.8377 2.2415 0.6830
FgAA (SFT) Llama3-8B 0.9218 0.1925 0.7813 0.7333 8.0528 9.9019 0.9698
FgAA (API) DeepSeek-V3.2-Exp 0.9118 0.1660 0.7573 0.6626 7.5210 9.8720 0.9960
Table 1: Main Baseline Comparison. Performance of RLAA against local anonymization methods (DP-BART+, IncogniText) and naive adversarial migration (FgAA). RLAA achieves the best privacy-utility trade-off, preventing the utility collapse observed in FgAA while avoiding the high hallucination of IncogniText.
Figure 3: Privacy-Utility trade-off of RLAA and FgAA in different iterations on two datasets.
Dataset: PersonalReddit
Original Text 1.0000 0.4442 1.0000 1.0000 10.0000 10.0000 1.0000
DeepSeek-V3.2-Exp Without Arb. 0.8264 0.2056 0.4649 0.2082 5.4380 9.3550 1.0000
RLAA 0.9240 0.2087 0.7401 0.6365 7.8350 9.9670 0.9920
Llama3-8B Without Arb. 0.7297 0.1948 0.2180 0.0533 3.5420 8.9330 0.9420
RLAA 0.8788 0.2130 0.5958 0.4251 7.0780 9.8090 0.9480
Qwen2.5-7B Without Arb. 0.8741 0.2302 0.6156 0.3973 6.8590 9.6120 0.9750
RLAA 0.9135 0.2531 0.7549 0.6413 7.8020 9.8620 0.9740
Dataset: reddit-self-disclosure
Original Text 1.0000 0.4943 1.0000 1.0000 10.0000 10.0000 1.0000
DeepSeek-V3.2-Exp Without Arb. 0.9118 0.1660 0.7573 0.6626 7.5210 9.8720 0.9960
RLAA 0.9132 0.1434 0.8070 0.7364 7.5620 9.9090 0.9920
Llama3-8B Without Arb. 0.8187 0.1591 0.4919 0.2805 5.3523 9.6629 0.9545
RLAA 0.8572 0.1136 0.6016 0.4703 6.8939 9.6932 0.9129
Qwen2.5-7B Without Arb. 0.8936 0.2000 0.6458 0.4800 7.2115 9.8269 0.9769
RLAA 0.9071 0.1753 0.7273 0.6208 7.7200 9.9320 0.9560
Table 2: Ablation Study and Generalization Test. The results demonstrate that the Arbitrator consistently improves the Privacy-Utility trade-off across different model scales.
Figure 4: MRS Analysis on Llama-3-8B. The marginal rate of substitution (MRS) represents the utility cost per unit of privacy gain. FgAA (red) shows a sustained increase (irrationality), while RLAA (blue) maintains a stable low MRS. Complete results for DeepSeek and Qwen are provided in Appendix D
4.2 Performance Analysis
Our evaluation consists of two parts: a comprehensive baseline comparison shown in Table 1 and an ablation study to verify the contribution of the rationality mechanism across different models shown in Table 2.
4.2.1 Baseline Comparison Results
Table 1 presents the comparative analysis on the PersonalReddit and reddit-self-disclosure datasets. RLAA demonstrates the most effective privacy-utility balance among local methods, achieving a high utility score of 0.8788 on PersonalReddit while preserving substantial structural integrity (ROUGE-L: 0.5958). In stark contrast, other local baselines suffer from severe trade-off imbalances: the naive adversarial migration (FgAA-Naive) experiences a catastrophic utility collapse (ROUGE-L: 0.2180) due to greedy over-editing, while IncogniText frikha2024incognitext sacrifices significant semantic structure (ROUGE-L: 0.3499) to achieve lower privacy risks via hallucination-prone attribute fabrication. Furthermore, our evaluation reveals the fragility of distillation-based methods in out-of-distribution scenarios. Despite utilizing advanced SFT and DPO alignment, SEAL kim2025self fails to prevent the utility collapse across both datasets (ROUGE-L of 0.1205 on PersonalReddit and 0.1213 on reddit-self-disclosure), which demonstrates that without rational framework design, distilled models still resort to destructive rewriting. Notably, RLAA’s architectural rationality effectively compensates for model scale. Despite running on a local 8B model, it remains highly competitive with the powerful remote baseline FgAA (API) (DeepSeek-V3.2-Exp), even surpassing it in Utility (0.8788 vs. 0.8264) on PersonalReddit while maintaining comparable privacy protection. The results confirms that a rational agent architecture can rival resource-intensive cloud services without the associated privacy paradox.
4.2.2 Ablation Study
To isolate the contribution of the proposed rationality mechanism, we treat FgAA as an ablation variant representing "Without Arb." Table 2 presents this ablation analysis across three model backbones to verify mechanism generalization. The comparison shown in Table 2 serves as a direct ablation of the arbitrator module. Removing the arbitrator consistently degrades utility across all models. On PersonalReddit with Llama3-8B, utility drops from 0.8788 to 0.7297. Crucially, the arbitrator also improves privacy protection. By filtering out noise, the system focuses edits on genuine leaks. On reddit-self-disclosure, RLAA achieves strict Pareto dominance, reducing privacy risk from 0.1591 to 0.1136. This confirms that the economic rationality constraint is not merely a safety brake but an active optimization component. Besides, We extend the ablation to model capabilities by evaluating on the SoTA level model DeepSeek-V3.2-Exp. Even with a powerful model less prone to hallucinations, RLAA still improves the utility score (+0.0976 on PersonalReddit) compared to the standard FgAA strategy. This serves as a stress test confirming that the irrationality issue is intrinsic to the greedy adversarial algorithm itself, not just a symptom of weak LSMs. RLAA functions as a general rationality layer that is effective regardless of the underlying model scale.
4.2.3 Economic Efficiency Analysis
To validate our theoretical premise regarding agent rationality, we analyze the marginal dynamics of the anonymization process. Figure 4 visualizes the Marginal Rate of Substitution (MRS) across iterations for Utility, ROUGE-L, and BLEU metrics.
As shown in Figure 4, the baseline FgAA (Red Curve) exhibits a continuous increase in MRS across all metrics and datasets, which represents a hyperinflation of the privacy price economically. This confirms that greedy agents inevitably drift into an irrational state. In contrast, RLAA (Blue Curve) maintains a low and stable MRS trajectory. By employing the arbitrator as a rationality gatekeeper, RLAA successfully identifies the point of diminishing returns and prevents the destructive over-anonymization. This ensures that the system operates efficiently along the Pareto Frontier. The comparison between the blue and red curves in Figure 4 serves as empirical proof of our core theorem: Constraint on rationality prevents utility collapse. RLAA not only merely fix bugs in LSMs but also fundamentally alters the economic behavior of the agent, transforming it from a greedy optimizer into a rational decision-maker.
5 Conclusion
In this paper, we address the critical "utility collapse" encountered when migrating adversarial anonymization to local small-scale models, and reframe it as a task of economic rationality. To resolve this, we propose RLAA (Rational Localized Adversarial Anonymization), a novel training-free framework featuring an Attacker-Arbitrator-Anonymizer architecture. RLAA integrates an arbitrator to explicitly validate adversarial reasoning and enforce a rational stopping criterion. This mechanism ensures economic efficiency in text modification, structurally preventing irrational over-anonymization driven by hallucinations and diminishing marginal returns. Extensive evaluations demonstrate that RLAA significantly outperforms our baselines, achieving a better privacy-utility trade-off and offering a viable path to resolving the privacy paradox.
6 Limitations
While RLAA demonstrates significant advancements, there are several limitations to acknowledge and to be mitigated in future work:
Firstly, the arbitrator module introduces additional inference steps, which requires an extra pass of reasoning validation. While this ensures economic efficiency, it increases the computational cost per iteration. We analyzed the detailed extra computational cost in Appendix B and proved that it is acceptable. Secondly, although we used DeepSeek-V3.2-Exp as a powerful adversary for evaluation, this only represents the boundary of the current large-scale model’s attack capability, and more advanced models may need to be evaluated in the future. Besides, unlike Differential Privacy (DP) methods, RLAA does not provide mathematically provable privacy budgets $\epsilon$, which represents a trade-off favoring semantic utility and naturalness over formal indistinguishability.
Appendix
Appendix A Theoretical Analysis
In this section, we provide a formal economic definition and proof to support our claim that naive adversarial strategies are inherently irrational, while RLAA is rational.
Definition 4 (Economic Rationality Condition.).
$\text{MRS}_{t}\leq\lambda$ (10)
A.1 Proof of RLAA’s Rationality
We hereby demonstrate how the proposed RLAA framework achieves structural rationality using discrete verification signals.
Assumption: Discrete Estimation.
While the marginal privacy gain $\Delta Pt$ is a continuous latent variable, LSM is unable to output accurate numerical estimates due to its limited capabilities and insensitivity to numbers. Therefore, we model the arbitrator’s discrete validity judgment $v\in\{\textsc{High},\textsc{Med},\textsc{Low},\textsc{Inv}\}$ as a quantized estimator of the expected privacy gain $\mathbb{E}[\Delta P]$. We define a mapping $\Phi$ that relates discrete labels to the expected gain:
where $\gamma$ represents a significant privacy improvement, and $\xi$ represents a negligible privacy gain (noise).
Proposition 1 (Rationality Guarantee).
RLAA ensures that the anonymization process remains within the rational zone where $\text{MRS}\leq\lambda$ by filtering operations where the expected denominator $\mathbb{E}[\Delta P]$ approaches zero.
Proof for Scenario 1: Hallucination Defense (Instantaneous Irrationality).
In a naive system, consider the attacker hallucinates a non-existent leak $l_{hall}$. Mathematically, since $l_{hall}$ is spurious, the privacy gain from modifying it is negligible ($\Delta P_{t}\to\xi\approx 0$). However, a naive anonymizer still executes a text modification to address it. Any text alteration inherently disrupts the original semantic structure, incurring a definitive utility cost $\Delta C_{t}\geq\epsilon>0$ (where $\epsilon$ is the minimum modification cost). Substituting these into Eq. 6 reveals a singularity:
$\lim_{\Delta P_{t}\to\xi}\text{MRS}_{t}=\frac{\Delta C_{t}}{\Delta P_{t}}\approx\frac{\epsilon}{\xi}\to\infty$ (12)
This represents an immediate deadweight loss. In contrast, RLAA employs the arbitrator to evaluate the reasoning logic $R$. Since $l_{hall}$ is a hallucination, the arbitrator assigns:
$v_{hall}=\mathcal{M}_{arb}(l_{hall},R)\in\{\textsc{Low},\textsc{Inv}\}$
Based on our mapping assumption (Eq. 11), this validity implies $\mathbb{E}[\Delta P]\approx\xi$. Consequently, the policy mapping function $\Pi_{select}$ triggers a rejection:
$\pi=\Pi_{select}(v_{hall})=\textsc{Ignore}$
Thus, the anonymizer performs no modification, ensuring the utility cost $\Delta C_{t}=0$. The transaction is canceled, maintaining rationality.
Proof for Scenario 2: Rational Equilibrium (Asymptotic Irrationality).
Let $\mathcal{L}^{(t)}$ be the set of leaks detected at iteration $t$. As $t$ increases, significant PIIs are resolved. The nature of remaining leaks shifts from explicit (e.g., names) to implicit (e.g., stylistic cues). For these long-tail leaks, the Marginal Privacy Gain decays asymptotically towards zero ($\lim_{t\to\infty}\Delta P_{t}=\xi$). However, modifications targeting these subtle cues still require structural alterations, sustaining a persistent utility cost ($\Delta C_{t}\geq\epsilon>0$). This causes the MRS of a naive agent to diverge:
$\lim_{t\to\infty}\text{MRS}_{t}=\frac{\Delta C_{t}}{\lim_{t\to\infty}\Delta P_{t}}\to\infty$ (13)
In RLAA, for all remaining $l\in\mathcal{L}^{(t)}$ at the tail, the arbitrator determines they lack sufficient evidence:
$\forall l\in\mathcal{L}^{(t)},\quad v_{l}\in\{\textsc{Low},\textsc{Inv}\}$
This results in an empty policy set $\mathcal{P}^{(t)}=\emptyset$. Algorithm 1 triggers the early stop condition. The process terminates at state $x^{(t)}$ rather than proceeding to an irrational state $x^{(t+1)}$. Thus, RLAA structurally converges to a rational equilibrium.
Table 3: Computational latency comparison of FgAA and RLAA on two datasets.
Dataset Method Avg. Latency Overhead
self-disclosure FgAA 14.7s / sample -
RLAA 21.6s / sample $47\%$
PersonalReddit FgAA 95.1s / sample -
RLAA 196.8s / sample $107\%$
Appendix B Computational Cost Analysis
To quantify the inference overhead introduced by the arbitrator module, we analyzed the average wall-clock latency per sample on NVIDIA RTX 5090 across two datasets. We maintained the same experimental configuration as described in Section 4.1, utilizing the Llama3-8B model.
As shown in Table 3, RLAA exhibits an adaptive computational cost that correlates with the difficulty of the privacy threat: On the reddit-self-disclosure dataset, which mostly contains direct or hallucinatory leaks, RLAA incurs a moderate latency overhead ($\approx 1.47\times$). On the PersonalReddit dataset, which involves multi-hop privacy leaks requiring deep reasoning, RLAA invests significantly more time ($\approx 2.07\times$ overhead). However, this additional computation is critical. As detailed in Section 4.2, the fast but greedy baseline (FgAA) suffers from utility collapse (ROUGE-L drops to 0.2180), while RLAA’s deliberate iterative process preserves structural integrity (ROUGE-L: 0.5958).
It is crucial to note that text anonymization is predominantly an offline or asynchronous workflow (e.g., sanitizing datasets before publication). In these quality-centric scenarios, latency is a secondary constraint compared to the structural integrity of the data. While RLAA introduces an inference overhead compared to greedy baselines, this cost is a strategic trade-off: it converts computational time into critical utility preservation, ensuring the sanitized data remains usable for downstream tasks.
Table 4: Hyperparameters for training-based baselines.
Method Learning Rate Batch Size Epoch Max Seq Length Specific Configs
FgAA (SFT) $1\text{e-}5$ $4\times 1$ 10 1024 $r=16,\alpha=32$
SEAL (SFT) $2\text{e-}4$ $4\times 2$ 1 4096 $r=16,\alpha=16$
SEAL (DPO) $5\text{e-}6$ $4\times 1$ 1 2048 $\beta=0.01,\alpha=16$
DP-BART $1\text{e-}5$ 16 50 300 $\epsilon=2500,\delta=10^{-6}$
@@ Appendix C Implementation Details
General Setup
All local LLM-based frameworks (RLAA, FgAA, IncogniText) employed Llama-3-8B and Qwen2.5-7B as base models. Models were loaded in half-precision (float16) on a single NVIDIA RTX 5090 GPU.
Table 4 details the hyperparameter settings for the training-based baselines. For FgAA (SFT), we performed standard supervised fine-tuning on the Llama-3-8B for 10 epochs to ensure convergence. In contrast, for the distillation-based SEAL, we strictly adhered to the official code release kim2025self, implementing its specific two-stage pipeline: an aggressive SFT phase (LR $2\text{e-}4$) followed by a conservative DPO phase (LR $5\text{e-}6$), both restricted to a single epoch to mitigate overfitting. All LLM fine-tuning utilized QLoRA (4-bit) for memory efficiency. The DP-BART-PR+ baseline was trained with a gradient clipping norm $C=5.0$ and a privacy budget $\epsilon=2500$.
Table 5 details the generation hyperparameters for each agent. In the RLAA framework, the arbitrator uses greedy decoding with $\tau=0$ to ensure deterministic validation. For IncogniText, we adopted a two-stage pipeline where the target persona is generated via the DeepSeek-V3.2-Exp API with a higher temperature to encourage diversity.
@@
Table 5: Generation Hyperparameters.
@@ Appendix D Detailed Experimental Results
Figure 5 presents the MRS analysis for DeepSeek-V3.2-Exp and Qwen2.5-7B, which exhibit consistent trends with Llama-3-8B reported in the main text. Figure 6 illustrates the training and evaluation loss curves for both the attacker and anonymizer components, fine-tuned on the Llama-3-8B. As shown in the figures, both models achieved stable convergence, confirming that the performance gaps observed in the main experiments are due to the inherent limitations of the naive adversarial strategy rather than insufficient training.
Figure 5: MRS Analysis for additional models.
Figure 6: SFT Training Dynamics. Loss curves for the Attacker (Top) and Anonymizer (Bottom) modules during the SFT migration process. The plots demonstrate healthy convergence patterns.

### v2

@@ Appendix B Inference Latency Analysis
11:   $\pi_{k}\leftarrow\Pi_{select}(v_{k})$
@@ Appendix C Implementation Details
This appendix provides implementation configurations, including the algorithmic procedure in Algorithm 1, computational environment, training recipes and generation hyperparameters.
Base Models & Environment. All local LLM-based frameworks (RLAA, FgAA, IncogniText) employed Llama-3-8B and Qwen2.5-7B as base models. To align with consumer-grade deployment scenarios, all models were loaded in half-precision (float16) on a single NVIDIA RTX 5090 GPU.
Training Configurations for Baselines. Table 4 details the specific hyperparameter settings for all training-based baselines.
  • FgAA (SFT): We performed standard supervised fine-tuning on Llama-3-8B for 10 epochs to ensure convergence. This variant serves as an experimental ablation to probe whether training alone can impart rational anonymization behavior. Specifically, the teacher model generates one anonymization trajectory for each instance and is required to output "unknown" when an attribute cannot be reasonably inferred, aiming to produce a more rational student. All LLM fine-tuning utilized QLoRA (4-bit) for memory efficiency.
  • SEAL: For this distillation-based baseline, we strictly adhered to the official implementation, including its two-stage pipeline: an SFT phase followed by a conservative DPO phase, executed as specified by the released recipe.
  • DP-BART-PR+: According to the official code, this baseline was trained with a gradient clipping norm $C=5.0$ and a privacy budget $\epsilon=2500$.
@@
Table 5: Specific Generation Hyperparameters. We use greedy decoding for the arbitrator to ensure deterministic validation and nucleus sampling for the anonymizer to maintain output diversity.
Generation Hyperparameters. Table 5 details the generation hyperparameters for each agent.
  • RLAA Settings: The arbitrator uses a temperature of 0 to ensure deterministic validity judgments, while the attacker and anonymizer use a slight temperature (0.1 and 0.5) to balance creativity and instruction following.
  • IncogniText Pipeline: We adopt a two-stage pipeline where the target persona is pre-generated offline via DeepSeek-V3.2-Exp to encourage stylistic diversity. Crucially, API usage is strictly limited to this one-time data preparation phase, ensuring the subsequent anonymization process remains fully localized.
Figure 5: Privacy-utility trade-offs via structural metrics (ROUGE-L and BLEU). Results are shown for PersonalReddit (Left) and reddit-self-disclosure (Right), demonstrating RLAA’s resistance to structural collapse.
@@ Appendix D Detailed Experimental Results
This section expands on the main analysis by providing structural evaluations, training stability checks and a quantitative economic analysis that reveals a counter-intuitive alignment paradox.
Structural Privacy-Utility Trade-offs. Figure 5 plots the trade-off dynamics for structural metrics (ROUGE-L and BLEU). Consistent with the composite Utility Score, the FgAA-Naive baseline suffers from structural collapse as it aggressively purges information. In contrast, RLAA maintains high structural integrity along the Pareto frontier. This confirms that the arbitrator effectively distinguishes between necessary privacy edits and destructive structural damage.
Figure 6: Cumulative MRS Profiles across Different Model Scales. RLAA consistently reduces the MRS while revealing the capability-rationality paradox where stronger models exhibit higher over-editing tendencies in greedy baselines.
Figure 7: Extended MRS Dynamics. These results across Utility, ROUGE-L, and BLEU metrics confirm the generalization of RLAA’s rational decision-making.
The Capability-Rationality Paradox: A Quantitative Analysis. We extend the economic analysis to Qwen2.5-7B and DeepSeek-V3.2-Exp, summarizing the comparative rationality profiles in Figure 6 and detailing their iterative dynamics in Figure 7. To quantify the impact of RLAA across different model capabilities, we calculate the Rationality Gain (percentage reduction of MRS) in Table 6. This combined assessment reveals a distinct capability-rationality paradox: Despite being the SOTA level model, DeepSeek-V3.2-Exp exhibits the highest rationality gain. As visualized in Figure 6 and quantified in Table 6, RLAA reduces DeepSeek’s MRS by a massive 66.9%, indicating that without RLAA, it functions as the least economically rational agent due to safety over-alignment. In contrast, Qwen2.5-7B shows the least rationality gain of 28.9%. Its greedy baseline (FgAA’s MRS of 1.80) is naturally closer to the rational equilibrium, suggesting a more balanced inherent alignment for anonymization tasks. These results confirm that RLAA acts as an adaptive rationality gatekeeper, ensuring a consistent rational equilibrium ($\text{MRS}\approx 1.1\text{-}1.7$) regardless of the base model’s inherent bias.
Metric DeepSeek Llama Qwen
FgAA’s MRS 3.32 3.80 1.80
RLAA’s MRS 1.10 1.74 1.28
Rationality Gain 66.9% 54.2% 28.9%
Table 6: Quantifying Rationality Correction. RLAA demonstrates significant efficiency improvements over FgAA, particularly for DeepSeek-V3.2-Exp.
Training Convergence Verification. To ensure the baseline failure is not due to under-fitting, Figure 8 illustrates the SFT training dynamics. Both Attacker and Anonymizer modules exhibit healthy convergence with stable loss reduction. This rules out optimization failure and confirms that the performance gap stems from the inherent economic irrationality of the greedy strategy itself.
Figure 8: SFT Training Dynamics. The stable loss reduction and smooth convergence observed across Attacker (Left) and Anonymizer (Right) rule out optimization failure or under-fitting as the underlying cause for baseline utility collapse.
Appendix E Human Evaluation
To empirically validate the superiority of our rationality mechanism and confirm the usability of the generated text, we conducted a three-way pairwise comparison. The evaluation was scaled to involve five independent and non-author annotators. Each annotator judged 50 randomized samples, yielding a total of 250 judgments. The results visualized in Figure 9 strongly confirm the effectiveness of our mechanism. RLAA achieved a dominant average win rate of 88.4% ($\sigma=5.9\%$), while the FgAA’s win rate was restricted to a mere 3.6%. This extreme disparity validates that FgAA’s greedy strategy on LSMs results in a systematic utility collapse, producing text largely deemed unusable by human readers. In contrast, RLAA’s decisive preference rating substantiates that the arbitrator effectively preserves semantic integrity. We also observed a strong consensus across all five reviewers, with a remarkably low standard deviation in win rates, which provides strong statistical evidence for inter-annotator consistency and ensures the objective reliability of these findings.
Figure 9: Human pairwise evaluation results. RLAA achieves a dominant win rate against FgAA-Naive while reflecting high inter-annotator consistency.
Appendix F Case Studies
We present three qualitative examples extracted from our human evaluation samples comparing RLAA against the FgAA baseline. These cases demonstrate how RLAA preserves semantic utility while effectively reducing privacy risks.
Case 1: Career & Location (WFH) [Original]: Ah yes, my sitting room now closely resembles NASA’s control centre with the amount of screens and tech there! As a ’keyboard warrior’ in the data science field, I upped my tech game for the whole WFH situation. Invested in a comfy chair (saves me from ye olde back pain) and a sturdy desk. Can’t put a price on comfort, ya know? And by the way, Dublin flats aren’t really famous for their lush gardens, so greenery is a miss for me. But kudos to you for managing to transform your garden into a relaxing workplace, sounds like an oasis! Tips on growing indoor greens on a windowsill though? My wee flat might benefit from a bit of green. Cheers!  [RLAA]: Ah yes, my sitting room now closely resembles a control centre with the amount of screens and tech there! As an individual, I’ve invested in a home office setup for the flexibility of remote work. Set up a dedicated workspace. It’s hard to put a price on comfort, isn’t it? And by the way, apartments aren’t really famous for their lush gardens, so greenery is a miss for me. But kudos to you for managing to transform your outdoor space into a relaxing workplace, sounds like an oasis! Tips on growing indoor greens on a windowsill though? My space might benefit from a bit of green. Cheers!  [FgAA]: Ah, yes, the space I’m in now closely resembles a space. Comfort is important, isn’t it? And by the way, I’m having trouble finding some plants. But I’d love some tips on growing plants.  Analysis: RLAA generalizes "NASA" $\to$ "control centre" and "Dublin" $\to$ "apartments", preserving the WFH narrative. FgAA strips almost all context.
Case 2: Profession & Specific City [Original]: oh yeah, i totally get what you mean. i’m also trying to shuffle things around my flat in Lisbon. last year, i decided to give some revamp to my workspace. ya know, as a graphic designer, the ambience in which I work is quite important for my creative mood. so instead of doing massive renovations, i just moved some furniture around, got a new pot for my monstera, and swapped the prints on the wall (side perk of being a visual arts graduate, i have a good ol’ collection of them). i also found some cool artsy stickers on etsy… didn’t burn a hole in my middle income wallet. up next, redo the bedroom. tryna stay within budget tho, lisbon life ain’t cheap ya know. any of you redditors got some cost-friendly DIY ideas? :D  [RLAA]: I’m totally getting what you mean. I’m also trying to refresh my space. Last year, I decided to give some revamp to a space. You see, the atmosphere in which I work is quite important for my productivity and focus. So instead of doing massive renovations, I just rearranged my furniture, got a new plant, and swapped the artwork on the wall (side perk of having a diverse selection of pieces)… didn’t break the bank. Next on my list is to refresh my bedroom. Being mindful of budgeting can be challenging, you know. Any of you got some affordable DIY ideas? :D  [FgAA]: I totally get what you mean. I’m trying to make some changes. The environment in which I work is quite important. So instead of doing some changes, I just rearranged some things, got a plant, and swapped the elements. The whole thing not only refreshed the space but also didn’t cost a lot. Got some ideas?  Analysis: RLAA neutralizes "Lisbon/Graphic Designer" while keeping the detailed renovation narrative. FgAA produces a repetitive, dry summary.
Case 3: Distinctive Voice & Persona [Original]: Mate, being an astronomer and all, I can tell ye I’ve spent hundreds, if not thousands, of nights under the stars. But hiking? I’d need a wee bit more physics for that, haha! And here in Edinburgh, you’re never too far from a cheeky hill to climb. Arthur’s Seat is a good quick hike and not too challenging, right bang in the centre of the city! You can see all the way to glasgow from the top on a clear day. Mind, it can get very nippy up there, this being Scotland and all… "There’s no such thing as bad weather, only the wrong clothes"!  [RLAA]: Mate, being someone who appreciates the beauty of the night sky, I can tell you I’ve enjoyed many moments of contemplation under the stars. But I’d need a bit more perspective for that, haha! And in a neighborhood, you’re never too far from a hill to climb. A nearby hill is a good quick hike and not too challenging, in a scenic area! You can see all the way to another nearby area from the top… Mind, it can get quite chilly up there… So there you go: "There’s no such thing as bad weather, only the wrong clothes"!  [FgAA]: Being someone who enjoys spending time outside, I can tell I’ve spent time in the outdoors. But venturing? I’d need more information before attempting it. And you’re never too far from a hill to climb. There’s a trail in the area! You can see a long way on a clear day. Mind, it can get quite cool up there. The hills though, now there’s a trail…  Analysis: RLAA retains the "Mate" persona and enthusiastic tone ("haha!", quotes), generalizing landmarks. FgAA completely loses the voice.